Circular No. 7741

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001fc IN UGC 11683 T. Puckett and L. Cox, Mountain Town, GA, report the discovery of

an apparent supernova (mag 17.5) on an unfiltered CCD frame (limiting mag 20.2) taken with the Puckett Observatory 0.60-m automated patrol telescope on Oct. 28.00 UT. SN 2001fc is located at α = 21^{h}08^{m}26^{s}.69, δ = +18^{o}11′17′′.0 (equinox 2000.0), which is 11′′.8 south of the center of UGC 11683. The new object was confirmed on CCD frames taken on Oct. 28.99. SN 2001fc is not present on images taken by Puckett on 2001 July 15 and 22 (limiting mag ∼ 20.5) or on Palomar Sky Survey images taken on 1992 Aug. 25 (limiting mag ∼ 21.0), 1990 Aug. 15 (limiting mag ∼ 21.0), and 1954 July 10 (limiting mag ∼ 20.0).

S/2001 (617) 1 W. J. Merline, Southwest Research Institute (SwRI); L. M. Close and

N. Siegler, University of Arizona; and D. Potter, University of Hawaii (UH), report for their collaboration (C. R. Chapman, SwRI; C. Dumas, Jet Propulsion Laboratory; F. Menard, Observatoire de Grenoble; D. C. Slater, SwRI) the discovery on Sept. 22.6 UT of a large companion to the Jupiter L_{5}-Trojan (617) Patroclus (predicted V = 15.1) from J-, H-, and K′-band direct imaging with the 8.1-m Gemini North Telescope (+ Hokupa’a adaptive optics system) on Mauna Kea. The two components are of similar size, differing in brightness by ∼ 0.2 mag. This is the first known binary among the Trojans. Confirming observations were acquired at the Gemini telescope on Oct. 13 by A. C. Baker and M. G. Edmunds (Cardiff), G. Mathlin (Bath), O. Guyon (UH), K. Roth (Gemini Observatory), and Merline. The data show the fainter component at the following separations and positions angles: Sept. 22.6247, 0′′.21, 251^{o}; Oct. 13.4828, 0′′.12, 219^{o}. In observations by Merline et al. on Sept. 21 and by Potter on Oct. 11, the binary was marginally resolved.

COMETS C/2001 T1 AND C/2001 T2 (SOHO) Further to IAUC 7730, D. Hammer reports his measurements for two

additional sungrazing comets found by X. Leprette and R. Kracht on SOHO website images. Full information appears on MPEC 2001-T52.

Comet  2001 UT  α_{2000}  δ_{2000}
C/2001 T1 Oct. 9.104  12^{h}54.^{m}5  − 4^{o}15′
C/2001 T2  12.154  12 50.1  − 8 50
2001 October 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green